OpenStudy (anonymous):

Since YouTube first became available to the public in mid-2005, the rate at which video has been uploaded to the site can be approximated by v(t) = 525,600(0.42t2 + 2.7t − 1) hours of video per year (0.5 ≤ t ≤ 7), where t is time in years since the start of 2005. (a) Find an expression for the total number of hours V(t) of video at time t (starting from zero hours of video at t = 0.5). V(t) = (b) Use the answer to part (a) to estimate the total number of hours of video uploaded by the start of 2010.

OpenStudy (anonymous):

For part a, isn't the answer the given function?

geerky42 (geerky42):

@ArkGoLucky None, we are given the rate of hours of video uploaded. What we want is total number of hours from beginning.

geerky42 (geerky42):

So since \(v(t)=V'(t)\) and we want to find \(V(t)\), we would need to take integral of \(v(t)\), from initial time \(t_o=0.5\) to desired time \(t\). Which would looks like \(\displaystyle\int_{0.5}^tv(x)~dx \\~\\\displaystyle= \int_{0.5}^t 525,600(0.42x^2 + 2.7x − 1)~dx\\~\\=\cdots~?\)

geerky42 (geerky42):

@rmwendt11